# Genetic Quantization-Aware Approximation for Non-Linear Operations in Transformers

### v1

@@ 1. Introduction
The advent of the Transformer-based neural networks marked a new era for natural language processing (kenton2019bert, ) and computer vision tasks (liu2021swin, ; xie2021segformer, ). The performance greatly benefits from the self-attention mechanism in Transformers, which could capture long-range dependencies well, but with a substantial overhead in computation and memory. Extensive research has been conducted to facilitate the deployment of Transformers on edge devices. Techniques like lightweight structure integrating convolution and linear attention (Cai_2023_ICCV, ; zhang2023augmented, ) emerge, while quantization (kim2021bert, ; pmlr-v202-liu23w, ) and run-time pruning (tu2023multcim, ) has become favored approaches to further reduced the hardware burden. However, the optimization of non-linear operations is frequently neglected in Transformer-based models which can be costly due to the frequent involvement of high-precision operations like 32-bit floating-point (FP32) or integer (INT32) arithmetic. As reported by (stevens2021softermax, ), the inefficiency of non-linear operations seriously impedes the speed-up of Transformers at lots of hardware platforms.
@@ 2.1. Non-Linear Operations in Transformers
Non-linear operations are ubiquitous in vanilla transformer models (vaswani2017attention, ). The multi-head self-attention employs Softmax to capture the correlation in different dimensions, while GELU acts as the activation function within the feed-forward network (FFN), and LayerNorm is for the normalization. However, as the demand for edge computing increases, numerous lightweight transformer architectures have emerged, such as Softmax-free linear attention (bolya2022hydra, ; han2023flatten, ) and hybrid models that incorporate depthwise-separable convolution into early stages or FFNs (Cai_2023_ICCV, ). The non-linear operations in lightweight transformers are diverse such as HSWISH and cosine. Generally, these non-linear operators require high-precision computations with disparate dataflow to ensure accuracy. Hence, it becomes imperative to explore methods for handling diverse non-linear operations within a unified and simple hardware engine.
@@ 4.3. Hardware Performance
In order to further prove the necessity and superiority of INT8 LUT-based $pwl$, we implement the two types of hardware units depicted in Figure (1) using Verilog HDL. The area and power dissipation of each LUT-based unit are obtained through synthesis using Synopsys Design Compiler, utilizing TSMC’s 28-nm Technology. To ensure a fair comparison, the operating frequency is set to 500MHz for all hardware units. The comparison results are shown in Table 6 where the FP32 denotes the LUT-based $pwl$ without input quantization representing the methods employed by NN-LUT and RI-LUT. The results reveal that an 8-entry INT8 LUT-based $pwl$ requires merely 961um2 of area, achieving remarkable reductions of 81.3% and 81.7% compared to high-precision FP32 and INT32 units, respectively. In terms of power dissipation, it consumes 0.4mW resulting in substantial savings of 80.2% in FP32 and 79.3% in INT32. Moreover, expanding the LUT storage to 16 entries results in an approximate 1.71$\times$ increase in area and 1.95$\times$ in power relative to the 8-entry INT8 configuration. Therefore, opting for a smaller entry size and low-precision LUT configuration proves to be crucial for $pwl$ under integer-only quantization.

### v2

@@ 1. Introduction
The advent of the Transformer-based neural networks marked a new era for natural language processing (kenton2019bert, ) and computer vision tasks (liu2021swin, ; xie2021segformer, ). The performance greatly benefits from the self-attention mechanism in Transformers, which could capture long-range dependencies well, but with a substantial overhead in computation and memory. Extensive research has been conducted to facilitate the deployment of Transformers on edge devices. Techniques like lightweight structure integrating convolution and linear attention (Cai_2023_ICCV, ; zhang2023augmented, ) emerge, while quantization (kim2021bert, ; pmlr-v202-liu23w, ; liu2023llm, ) and run-time pruning (tu2023multcim, ) has become favored approaches to further reduced the hardware burden. However, the optimization of non-linear operations is frequently neglected in Transformer-based models which can be costly due to the frequent involvement of high-precision operations like 32-bit floating-point (FP32) or integer (INT32) arithmetic. As reported by (stevens2021softermax, ), the inefficiency of non-linear operations seriously impedes the speed-up of Transformers at lots of hardware platforms.
@@ 2.1. Non-Linear Operations in Transformers
Non-linear operations are ubiquitous in vanilla transformer models (vaswani2017attention, ). The multi-head self-attention employs Softmax to capture the correlation in different dimensions, while GELU acts as the activation function within the feed-forward network (FFN), and LayerNorm is for the normalization. However, as the demand for edge computing increases, numerous lightweight transformer architectures have emerged, such as Softmax-free linear attention (han2023flatten, ) and hybrid models that incorporate depthwise-separable convolution into early stages or FFNs (Cai_2023_ICCV, ). The non-linear operations in lightweight transformers are diverse such as HSWISH and cosine. Generally, these non-linear operators require high-precision computations with disparate dataflow to ensure accuracy. Hence, it becomes imperative to explore methods for handling diverse non-linear operations within a unified and simple hardware engine.
@@ 4.3. Hardware Performance
In order to further prove the necessity and superiority of INT8 LUT-based $pwl$, we implement the two types of hardware units depicted in Figure (1) using Verilog HDL. The area and power dissipation of each LUT-based unit are obtained through synthesis using Synopsys Design Compiler, utilizing TSMC’s 28-nm Technology. To ensure a fair comparison, the operating frequency is set to 500MHz for all hardware units. The comparison results are shown in Table 6 where the FP32 denotes the LUT-based $pwl$ without input quantization representing the methods employed by NN-LUT and RI-LUT. The results reveal that an 8-entry INT8 LUT-based $pwl$ requires merely 961um2 of area, achieving remarkable reductions of 81.3% and 81.7% compared to high-precision FP32 and INT32 units, respectively. In terms of power dissipation, it consumes 0.4mW resulting in substantial savings of 80.2% in FP32 and 79.3% in INT32. Moreover, expanding the LUT storage to 16 entries results in an approximate 1.71$\times$ increase in area and 1.95$\times$ in power relative to the 8-entry INT8 configuration. Therefore, a smaller entry size and low-precision are crucial for $pwl$ under integer-only quantization.